# The PAU Survey: galaxy stellar population properties estimate with narrow-band data

### v1

@@ 5.3 Comparison with COSMOS2020
Figure 12 depicts the overall distribution contours between the COSMOS2020 data, and our cigale and Prospector estimates with narrow-band PAUS and broad-band CFHTLS observations on the matched sample. We find that the properties estimated with cigale match more closely those from the COSMOS2020 catalogue, in particular for the stellar mass and the stellar ages, while Prospector better matches the SFR distribution of COSMOS2020. Even though COSMOS2020 finds a larger population of red sequence objects towards lower SFRs, the SFMS is mostly similarly recovered with both of our codes and within COSMOS2020. Ages on the other hand show alike distributions between COSMOS2020 and cigale, however with a offset of around 1 Gyr, where either COSMOS2020 underestimates the ages, or cigale overestimates them. As previously shown, the Prospector results on ages strongly differ from the other two samples, which we relate to the aforementioned different SFH treatment that allows overall higher ages with a more pronounced right tail of the distribution.
@@
5.4 Differences to previous narrow-band results
Previous studies of SED fitting with narrow-band data have been performed as part of the miniJPAS and J-PAS surveys, where the effect of an increase in photometric resolution and of a change of fitting code on the SED and physical property estimates was analyzed (Mejía-Narváez et al., 2017; González Delgado et al., 2021). These studies found reduced biases in the physical parameter estimation with miniJPAS narrow-bands with respect to broad-band data. Moreover, they determined that the age-metallicity degeneracy is mostly hidden at broad-band resolutions, only emerging at quasi-spectroscopic resolutions. While we cannot confirm a reduction in deviations between narrow-bands and broad-bands on these parameters, the larger parameter space indicates some degree of improvement due to the higher resolution.
@@ 5.4 Differences to previous narrow-band results
Overall, such discrepancies between similar studies again emphasize how unstable physical property estimates even at similar resolutions can be. If the SPS model assumptions, stellar spectral libraries, or inferences frameworks differ, varying biases in the recovery of physical properties of stellar populations of galaxies have to be expected. This arises especially for ages, dust attenuation parameters and the mean stellar metallicity, given the difficulty in their estimation with clearly resolved emission lines or other spectral features. Recent work by Nersesian et al. (2024) also concluded that stellar populations properties from SED fitting with Prospector are dominated by the modelling approach even with deep photometric data with a broad baseline, including narrow-bands. This, combined with the differences between our results and previous narrow-band studies, indicates that the increased spectral resolution of current narrow-band surveys does not yield an improved robustness of galaxy physical properties measurements, due to the large biases induces by the SPS modelling choices.
@@ 6 Conclusion
In this work, we tested SED fitting to quantify the effect of optical narrow-band photometries from the PAU Survey on the predictive power of galaxy stellar population properties with modern SPS codes.
@@
We find that while the accessible range of parameter values is increased with the narrow-band filters using cigale, the overall difference in comparison with $ugriz$ data only is not strongly pronounced. This can be caused by multiple reasons: non-flexible SPS models, inefficient parameter space coverage, likelihood maximization difficulties or low narrow-band signal-to-noise ratio. With Prospector, tests have to be performed to validate whether the metallicity issue might be caused by the rise of the age-metallicity degeneracy with increased spectral resolution, for example by comparing to broad-band fitting only.
@@
Still, we find that the PAUS filter set, as well as the CFHTLS broad-bands, both of which only cover the optical range and thus cannot rely on for example UV or IR bands for a more robust estimation of related properties, such as the SFR, are still able to recover previous measurements with larger baselines, that is COSMOS2020. With the majority of our sample covering a redshift range where the 4000 Å break lies within the regime of the PAUS bands, a more reliable SFH modelling is facilitated that allows robust SFR constraints.
@@
Furthermore, we show that the Prospector-$\alpha$ model is able to recover stellar masses and SFRs of the COSMOS sample, although with similar values compared to cigale. This code provides moreover increased flexibility in its age range and can provide SEDs with more diverse dust and metallicity values. To test how much of this discrepancy compared to cigale and COSMOS2020 depends on the PAUS filters will have to be further explored by Prospector fits on only $ugriz$ broad-bands. As this however requires an extreme amount of computing resources for MCMC chain convergence per galaxy, the applicability to large data sets is heavily constrained. While we expect the Prospector code to be able to create more diverse SEDs with physically motivated SPS models, the broader range of age-dust-metallicity values could be either related to superior SED forward-modelling, or simply be caused by incomplete coverage of the prior space due to the limited number of iterations.
@@
Differences between the results from both codes can occur for a number of reasons. The statistical framework, prior choices, different models for contributions to the SED, as well as the choice of stellar spectral templates all influence the quality of the best-fit spectrum and its derived physical properties. While these do not seem to have a large effect on stellar masses and even SFRs, the differences in the age-dust-metallicity space indicate that SED fitting results should not be over-interpreted in galaxy evolution studies, even with large baselines or quasi-spectroscopic resolutions. This is mainly due to the property degeneracies, but also caused by model and template choices in the applied SED code, as well as its statistical inference method.
@@
Overall, we find that while the PAUS narrow-band photometry provides a large improvement on the accuracy of photometric redshifts, the effect on the estimation of physical properties of galaxy stellar populations is not very significant. With stellar masses and SFRs being mostly derived robustly with broad-bands only, slight improvements can in general be expected for ages, metallicities or dust attenuation model parameters. Still, the choice of SED fitting code, stellar spectral library as well as for example dust or gas models have a large impact too. This thus requires caution in the interpretation of physical parameters from photometry in galaxy studies and shows the clear drawbacks of even well resolved photometry in comparison the spectroscopy.
@@
Finally, we obtain a catalogue of new age, dust and metallicity measurements on the PAUS-COSMOS sample, that provides a more flexible and physically meaningful estimation, as it does not rely on simplified template fitting or fixing multiple SPS parameters (such as $Z_{*}$ or the dust attenuation curve) to reduce the dimensionality of the prior.
@@ Appendix A Prospector example posterior distributions
Figure 14 shows an example of a fit from Prospector where the chains do not converge, with many properties showing posteriors with multiple peaks. This is not due to short sampling, as the shown plot was obtained after MCMC with 64 walkers for 8192 iterations each, leading to overall >500 000absent500000>500\,000> 500 000 likelihood calls. The stellar mass is seemingly unaffected, with a tight constraint, but metallicity, the slope modifier of the dust attenuation model and some SFR ratios clearly show a degenerate behaviour. While this is not apparent for all galaxies, it still showcases how the correlation between dust, metallicity and SFH can account for the large differences between simulated ground truth and fitted properties. Moreover, it proves that simply sampling longer does not necessarily improve the results as the degeneracy is not easily broken.
Figure 14: Posterior distribution of an example galaxy with non-converged properties from Prospector. The properties shown are the metallicity, the diffuse dust optical depths $\hat{\tau}_{\lambda}$, the ratio between diffuse dust and birth cloud optical depths, the dust slope modifier $n$, the stellar mass and the ratios of SFRs between adjacent age bins, which define the non-parametric SFH.
We next also illustrate how an increase of likelihood calls can positively affect the (or change in general) the posterior distribution and thus the values of the best-fit SED model. Figure 15 depicts the results from two Prospector runs on the same object, once with our nominal setup and once using 64 walkers with 8192 iteration each. While the stellar mass, as well as the SFR bin ratios of the non-parametric SFH do mostly agree within their $1\sigma$ contours, the ratio of diffuse and birth cloud dust optical depth and the stellar metallicity can shift significantly given more likelihood calls. This shows how the metallicity and dust property estimates can be unstable during the sampling process and show no clear convergence. However, this does not necessarily result in better results after longer sampling, due to the dust-age-metallicity degeneracy, where multiple points in this parameter space might have similar high likelihoods. Moreover, the shown example is not fully representative, as it is just a random sample. This can and will look significantly different depending on the properties of the objects that is fitted. Still, it shows that we can have confidence in our stellar mass and SFR estimates, but caution is required for galaxy evolution studies in general when relying on SED fitting to infer specific physical parameters such as the mean stellar ages or metallicities.
@@ Appendix B Comparison of SED with varying SSPs
As aforementioned, the SSP templates available in cigale and Prospector differ, leading to biases in the SED modelling given the same set of parameters. We here show in Fig. 16 a comparison plot between the output simulated composite stellar populations given the same set of input physical properties. The simulations assume a fixed metallicity $Z=Z_{\odot}$, a redshift $z=0.0$, and a delayed-$\tau$ shape of the SFH with $\langle t_{*}\rangle=5\,$Gyr, $\tau_{*}=2\,$Gyr, with no dust attenuation, dust emission or nebular emission. Thus, the models only differ by their applied stellar templates, with BC03 for cigale and MILES for Prospector. The first subplot shows the overall rest-frame SEDs derived from both codes for the input parameter set, with the second subplot depicting the residual between the two curves. Below, we plot the magnitude differences for the observed photometry in the PAUS and CFHTLS bands extracted from the observed-frame spectra. Therein, we see that the MILES templates produce overall brighter magnitudes than BC03. The relative magnitude difference decreases with the wavelength up to the $g$-band, but increases again towards to near-infrared. Such differences in the UV, optical and infrared thus naturally lead to differences in the physical parameters, if the same photometry is fit using varying SSPs.

### v2

@@ 5.3 Comparison with COSMOS2020
Figure 12 depicts the overall distribution contours between the COSMOS2020 data, and our cigale and Prospector estimates with narrowband PAUS and broadband CFHTLS observations on the matched sample. We find that the properties estimated with cigale match more closely those from the COSMOS2020 catalogue, in particular for the stellar mass and the stellar ages, while Prospector better matches the SFR distribution of COSMOS2020. Even though COSMOS2020 finds a larger population of red sequence objects towards lower SFRs, the SFMS is mostly similarly recovered with both of our codes and within COSMOS2020. The ages, on the other hand, show similar distributions between COSMOS2020 and cigale; however, with a offset of around 1 Gyr, where either COSMOS2020 underestimates the ages or cigale overestimates them. As previously shown, the Prospector results on ages strongly differ from the other two samples, which we relate to the aforementioned different SFH treatment that allows overall higher ages with a more pronounced right tail of the distribution.
@@
5.4 Differences to previous narrowband results
Previous studies of SED fitting with narrowband data have been performed as part of the miniJPAS and J-PAS surveys, where the effect of an increase in photometric resolution and of a change of fitting code on the SED and physical property estimates was analyzed (Mejía-Narváez et al., 2017; González Delgado et al., 2021). These studies found reduced biases in the physical parameter estimation with miniJPAS narrowbands with respect to broadband data. Moreover, they determined that the age-metallicity degeneracy is mostly hidden at broadband resolutions, only emerging at quasi-spectroscopic resolutions. While we cannot confirm a reduction in deviations between narrowbands and broadbands on these parameters, the larger parameter space indicates some degree of improvement due to the higher resolution.
@@ 5.4 Differences to previous narrowband results
Overall, such discrepancies between similar studies again emphasize how unstable physical property estimates even at similar resolutions can be. If the SPS model assumptions, stellar spectral libraries, or inferences frameworks differ, varying biases in the recovery of physical properties of stellar populations of galaxies have to be expected. This arises especially for ages, dust attenuation parameters and the mean stellar metallicity, given the difficulty in their estimation with clearly resolved emission lines or other spectral features. Recent work by Nersesian et al. (2024) also concluded that stellar populations properties from SED fitting with Prospector are dominated by the modeling approach even with deep photometric data with a broad baseline, including narrowbands. This, combined with the differences between our results and previous narrowband studies, indicates that the increased spectral resolution of current narrowband surveys does not yield an improved robustness of galaxy physical properties measurements, due to the large biases induces by the SPS modeling choices.
@@ 6 Conclusion
In this work, we tested SED fitting to quantify the effect of optical narrowband photometries from the PAU Survey on the predictive power of galaxy stellar population properties with modern SPS codes. We find that while the accessible range of parameter values is increased with the narrowband filters using cigale, the overall difference in comparison with $ugriz$ data only is not especially pronounced. This can be caused by multiple reasons: non-flexible SPS models, inefficient parameter space coverage, likelihood maximization difficulties, or low narrowband S/N values. With Prospector, tests have to be performed to validate whether the metallicity issue might be caused by the rise of the age-metallicity degeneracy with increased spectral resolution, for example by comparing to broadband fitting only.
@@
Still, we find that the PAUS filter set, as well as the CFHTLS broadbands, both of which only cover the optical range and thus cannot rely on for example UV or IR bands for a more robust estimation of related properties, such as the SFR, are still able to recover previous measurements with larger baselines, that is COSMOS2020. With the majority of our sample covering a redshift range where the 4000 Å break lies within the regime of the PAUS bands, a more reliable SFH modeling is facilitated that allows robust SFR constraints.
@@
Furthermore, we show that the Prospector-$\alpha$ model is able to recover stellar masses and SFRs of the COSMOS sample, although with similar values compared to cigale. This code provides moreover increased flexibility in its age range and can provide SEDs with more diverse dust and metallicity values. To test how much of this discrepancy compared to cigale and COSMOS2020 depends on the PAUS filters will have to be further explored by Prospector fits on only $ugriz$ broadbands. As this however requires an extreme amount of computing resources for MCMC chain convergence per galaxy, the applicability to large data sets is heavily constrained. While we expect the Prospector code to be able to create more diverse SEDs with physically motivated SPS models, the broader range of age-dust-metallicity values could be either related to superior SED forward-modeling, or simply be caused by incomplete coverage of the prior space due to the limited number of iterations.
@@
Differences between the results from both codes can occur for a number of reasons. The statistical framework, prior choices, different models for contributions to the SED, as well as the choice of stellar spectral templates all influence the quality of the best-fit spectrum and its derived physical properties. While these do not seem to have a large effect on stellar masses (or even SFRs), the differences in the age-dust-metallicity space indicate that SED fitting results should not be over-interpreted in galaxy evolution studies, even with large baselines or quasi-spectroscopic resolutions. This is mainly due to the property degeneracies, but also caused by model and template choices in the applied SED code, as well as its statistical inference method.
@@
Overall, we find that while the PAUS narrowband photometry provides a large improvement on the accuracy of photometric redshifts, the effect on the estimation of physical properties of galaxy stellar populations is not very significant. With stellar masses and SFRs being mostly derived robustly with broadbands only, slight improvements can in general be expected for ages, metallicities, or dust-attenuation model parameters. Still, the choice of SED fitting code, stellar spectral library as well as for example dust or gas models have a large impact too. Thus, this approach requires caution in the interpretation of physical parameters from photometry in galaxy studies and shows the clear drawbacks of even well resolved photometry in comparison the spectroscopy.
@@
Finally, we have obtained a catalogue of new age, dust and metallicity measurements on the PAUS-COSMOS sample, which provides a more flexible and physically meaningful estimation, as it does not rely on simplified template fitting or fixing multiple SPS parameters (e.g., $Z_{*}$ or the dust attenuation curve) to reduce the dimensionality of the prior.
@@ Appendix A Prospector example posterior distributions
Figure 14 shows an example of a fit from Prospector where the chains do not converge, with many properties showing posteriors with multiple peaks. This is not due to short sampling, as the shown plot was obtained after MCMC with 64 walkers for 8192 iterations each, leading to an overall outcome of >500 000absent500000>500\,000> 500 000 likelihood calls. The stellar mass is seemingly unaffected, with a tight constraint, but metallicity, the slope modifier of the dust attenuation model and some SFR ratios clearly show a degenerate behavior. While this is not apparent for all galaxies, it still showcases how the correlation between dust, metallicity, and SFH can account for the large differences between simulated ground truth and fitted properties. Moreover, it proves that simply sampling longer does not necessarily improve the results as the degeneracy is not easily broken.
Figure 14: Posterior distribution of an example galaxy with non-converged properties from Prospector. The properties shown are the metallicity, the diffuse dust optical depths $\hat{\tau}_{\lambda}$, the ratio between diffuse dust and birth cloud optical depths, the dust slope modifier, $n$, the stellar mass, and the ratios of SFRs between adjacent age bins, which define the non-parametric SFH.
Next, we illustrate how an increase of likelihood calls can positively affect the (or change in general) the posterior distribution and thus the values of the best-fit SED model. Figure 15 depicts the results from two Prospector runs on the same object, once with our nominal setup and once using 64 walkers with 8192 iteration each. While the stellar mass, as well as the SFR bin ratios of the non-parametric SFH do mostly agree within their $1\sigma$ contours, the ratio of diffuse and birth cloud dust optical depth and the stellar metallicity can shift significantly given more likelihood calls. This shows how the metallicity and dust property estimates can be unstable during the sampling process and show no clear convergence. However, this does not necessarily result in better results after longer sampling, due to the dust-age-metallicity degeneracy, where multiple points in this parameter space might have similar high likelihoods. Moreover, the shown example is not fully representative, as it is just a random sample. This can and will look significantly different, depending on the properties of the objects that is fitted. Still, it shows that we can have confidence in our stellar mass and SFR estimates, but caution is required for galaxy evolution studies in general when relying on SED fitting to infer specific physical parameters such as the mean stellar ages or metallicities.
@@ Appendix B Comparison of SED with varying SSPs
As aforementioned, the SSP templates available in cigale and Prospector differ, leading to biases in the SED modeling given the same set of parameters. We here show in Fig. 16 a comparison plot between the output simulated composite stellar populations given the same set of input physical properties. The simulations assume a fixed metallicity $Z=Z_{\odot}$, a redshift $z=0.0$, and a delayed-$\tau$ shape of the SFH with $\langle t_{*}\rangle=5\,$Gyr, $\tau_{*}=2\,$Gyr, with no dust attenuation, dust emission or nebular emission. Thus, the models only differ by their applied stellar templates, with BC03 for cigale and MILES for Prospector. The first subplot shows the overall rest-frame SEDs derived from both codes for the input parameter set, with the second subplot depicting the residual between the two curves. Below, we plot the magnitude differences for the observed photometry in the PAUS and CFHTLS bands extracted from the observed-frame spectra. Therein, we see that the MILES templates produce overall brighter magnitudes than BC03. The relative magnitude difference decreases with the wavelength up to the $g$-band, but increases again towards to near-infrared. Such differences in the UV, optical and infrared thus naturally lead to differences in the physical parameters, if the same photometry is fit using varying SSPs.